KpopAsianDramaFanYT:

Determine whether the function shown in the graph is even or odd. the graph will be down in the reply Answers: The function is even because it is symmetric with respect to the y-axis The function is odd because it is symmetric with respect to the y-axis. The function is even because it is symmetric with respect to the origin. The function is odd because it is symmetric with respect to the origin.

KpopAsianDramaFanYT:

1 attachment
aquire:

If a function only cares about the quantity of the input, not if it's positive or negative then that function is an even function.

snowflake0531:

The main thing to remember when looking at a graph and figuring whether it is odd or even, is that even functions are symmetrical with the y-axis, while for odd functions, if you spin the graph 180 degrees, looking at it upside-down, if it is still the same, it is an odd function

surjithayer:

if f(-x)=f(x),then it is even. if f(-x)=-f(x) then it is even.

aquire:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @surjithayer if f(-x)=f(x),then it is even. if f(-x)=-f(x) then it is even. \(\color{#0cbb34}{\text{End of Quote}}\) I think what you meat to say in the the second line is odd.

surjithayer:

YES it is odd,typing mistake.